### v2

# Converging to Stability in Two-Sided Bandits: The Case of Unknown Preferences on Both Sides of a Matching Market

Abstract.
We study the problem of repeated two-sided matching with uncertain preferences (two-sided bandits), and no explicit communication between agents. Recent work has developed algorithms that converge to stable matchings when one side (the proposers or agents) must learn their preferences, but the preferences of the other side (the proposees or arms) are common knowledge, and the matching mechanism uses simultaneous proposals at each round. We develop new algorithms that provably converge to stable matchings for two more challenging settings: one where the arm preferences are no longer common knowledge, and a second, more general one where the arms are also uncertain about their preferences. In our algorithms, agents start with optimistic beliefs about arms’ preferences and update these preferences over time. The key insight is in how to combine these beliefs about arm preferences with beliefs about the value of matching with an arm conditional on one’s proposal being accepted when choosing whom to propose to.
conference: ; ;
1. Introduction
The classic literature on two-sided matching (Gale and Shapley, 1962; Roth and Xing, 1997; Haeringer and Wooders, 2011, e.g.), encompassing applications including long- and short-term labor markets, dating and marriage, school choice, and more, has typically focused on situations where agents are aware of their own preferences. The problem of learning preferences while participating in a repeated matching market first started receiving attention in the AI literature in the work of Das and Kamenica (2005), and the general idea of two-sided matching under unknown preferences has since been studied in economics and operations research as well  (Lee and Schwarz, 2009; Johari et al., 2022). This area of research has received renewed attention in the last few years, along with novel theoretical insights into convergence properties of upper-confidence-bound (UCB) style algorithms (Liu et al., 2021; Kong et al., 2022; Zhang et al., 2022).
The two-sided matching problem involves agents on two sides of a market who have preferences for each other but cannot communicate explicitly. The goal is to create a matching process that ensures stability, where no pairs of agents would rather be matched with each other over their current match. The existence of such matchings was famously demonstrated constructively in the Gale-Shapley algorithm (Gale and Shapley, 1962) which structured around one side of the market proposing and the other side choosing whether to accept proposals. This theory has been applied to various markets, such as matching medical students to residencies (Roth and Peranson, 1999) and students to schools (Abdulkadiroğlu et al., 2005), with the assumption that agents know their own preferences. Considerable interest has also been shown in the AI community regarding two-sided matching in the presence of various constraints, such as diversity constraints (Aziz et al., 2022; Benabbou et al., 2018). A version of the problem, called two-sided bandits, was introduced in Das and Kamenica (2005), where agents engage in repeated matching without prior knowledge of preferences, and the impact of the matching mechanism on convergence to stability was studied. Recent work has also focused on computing the probability that a matching is stable or finding the matching with the highest probability of being stable under different models of preference uncertainty (Aziz et al., 2020).
The exploration-exploitation dilemma that characterizes bandit problems is further complicated in two-sided bandits by uncertainty on not just the values of the “arms” of the bandit, but also uncertainty on whether the arms will accept or reject your attempt to pull them (your proposal). This additional uncertainty arises from the multi-agent nature of the problem; instead of there being just one player, there are multiple players competing for the arms, and the arms also have preferences associated with the players. A popular solution choice for this type of problem involves the explore-then-commit style of algorithms (Pagare and Ghosh, 2023; Zhang and Fang, 2024). This flavor of solution includes a first phase of “learning” preferences, and a second phase using what has been learned in the prior one to commit to a choice. This involves spending some dedicated time trying to figure out the preferences of the agents. In this paper we are more interested in approaches where learning happens simultaneously with the matching process. One such algorithm, CA-UCB (Liu et al., 2021), assumes that at each round all the agents on the receiving side (henceforth arms) get to view all their proposals before deciding which one to accept. The algorithm provably converges to stability under the assumptions that (1) all arms have complete knowledge of their preferences, and (2) these preferences are common knowledge, so the proposers (henceforth players) also know them. The algorithm is a variation of the well-known family of UCB algorithms.
Contributions
In this paper, we are interested in the more general case, where neither the players nor the arms start off with knowledge of their own preferences. This problem is significantly more complex because it adds a layer of uncertainty that could potentially lead to agents converging to an unstable equilibrium more easily than in the case with one-sided uncertainty and common knowledge of the other side’s preferences. We tackle this in two steps. First, we relax the common knowledge assumption and assume only that arms are aware of their own preferences.
Recent works have used frameworks where players communicate with each other to find a matching in this setting (Zhang et al., 2022; Zhang and Fang, 2024). While there are approaches which allow for and design explicit communication frameworks (Zhang et al., 2022), and also others where players only communicate indirectly to the extent of carefully orchestrating “collisions” (Zhang and Fang, 2024), we resolve the question of convergence in the more traditional non-cooperative setting of stable matching when arm preferences are not common knowledge. Our algorithm uses the concept of “plausible sets” (arms that an player might be able to “win” in the next time step) from the CA-UCB algorithm (Liu et al., 2021). We show how to define and update these sets in order to develop a provably convergent algorithm (which we call Optimistic CA-UCB or OCA-UCB). The case with two-sided uncertainty is more challenging. We introduce the Probabilistic Conflict Avoiding - Simultaneous Choice Algorithm (PCA-SCA), a provably convergent framework that leverages an additional optimistic estimate for the probability of winning each contentious situation. Although we prove convergence using UCB-style reward estimates, the framework accommodates other methods such as Thompson Sampling. Experimentally, we demonstrate convergence to stability under both UCB and Thompson-based reward estimates and evaluate how the convergence speed scales with the number of agents and the degree of preference heterogeneity.
2. Setting
There are $N$ proposers (henceforth players) and $K$ proposees (henceforth arms). To ensure that each player can be matched, we assume that $N\leq K$ (Liu et al., 2020, 2021; Basu et al., 2021; Sankararaman et al., 2021). The set of players is denoted by $\left\{p_{i}\right\}_{i=1}^{N}$ and the set of arms by $\left\{a_{k}\right\}_{k=1}^{K}$. For each $p_{i}$, there is a distinct (unknown) mean reward $\mu_{i}^{k}$ associated with each arm $a_{k}$ with unit variance. If $\mu_{i}^{k}>\mu_{i}^{j}$, we say that $p_{i}$ prefers $a_{k}$ over $a_{j}$ denoted by $a_{k}\succ_{p_{i}}a_{j}$. Similarly, the arms also have distinct mean rewards associated with each player. The arms may or may not have this information available to them a priori depending on the setting in consideration.
At each time-step $t\in\{1,\cdots,T\}$, player $p_{i}$ attempts to pull an arm $a_{k}$. Let $A_{i}(t)$ and $\bar{A}_{i}(t)$ denote, respectively, the arm that player $p_{i}$ attempted to pull, and the arm that player $p_{i}$ successfully pulled at time $t$. If multiple players attempt to pull the same arm, a conflict will arise. If the arms are aware of their preferences, the player that is the most preferred by the arm will be picked in the event of a conflict (if the arms are not fully aware of their preferences, then the chosen player will depend on the arm’s decision-making algorithm). At the end of each time-step, if a player $p_{i}$ is successfully matched with an arm i.e. $\bar{A}_{i}(t)=m_{i}(t)$, it will receive a stochastic reward $X_{i}^{m_{i}(t)}\sim\mathcal{N}(\mu_{i}^{m_{i}(t)},1)$. Players that fail to successfully pull an arm will receive a reward of 0 and $\bar{A}_{i}(t)=\emptyset$. The final matching is made public to all the agents at the end of each time step.
We use the classic notion of a stable matching: a bipartite matching from the group of players to the group of arms such that no player and arm prefer each other over their current matching i.e $\nexists(p_{i},m_{i}(t))$ such that $p_{i}$ prefers an arm $a_{j}\neq m_{i}(t)$ and $a_{j}$ also prefers $p_{i}$ over its current match. There may be more than one stable match. Our primary focus in this work is on convergence to stability. Another quantity we study is maximum Player Pessimal Regret. We use the Gale-Shapley algorithm with arms as proposers in order to calculate the player-pessimal stable match (Gale and Shapley, 1962). Denote the reward received by player $p_{i}$ in a player pessimal match by $X_{i}$. Then, we define maximum player-regret at each time-step as $R(t)=\underset{i\in N}{\max}\left[X_{i}-X_{i}^{m(t)}\right]$ which is the maximum difference (over players) between the reward a player received in its current match and the reward it would have received in a stable matching that player-pessimal.
In this work we consider matching markets with varying levels of information availability. As is standard in models of learning in two-sided matching, we assume no explicit communication between players on the proposing side. Throughout the rest of the paper we assume that the players do not have access to their own preferences and need to learn them over time. As mentioned above, we consider several scenarios in terms of arm’s preferences. For convenience, we name those as follows:
  • Scenario 0: Players have access to arm preferences. Arms have access to arm preferences. We call this the Arm Preferences Common Knowledge (APCK) model.
  • Scenario 1: Players do not have access to arm preferences. Arms have access to arm preferences. We call this the Arm Preferences Known but Private (APKP) model.
  • Scenario 2: Players do not have access to arm preferences. Arms do not have access to arm preferences. We call this the Arm Preferences Unknown (APU) model.
The APCK model has been studied and a decentralized solution approach is described in the Conflict Avoiding Upper Confidence Bound (CA-UCB) algorithm (Liu et al., 2021). Our main contribution is to detail decentralized approaches to solutions for Scenarios 1 and 2. We begin with a quick recap of CA-UCB in the APCK model.
3. APCK Model and CA-UCB
In a recent paper, the APCK model (Scenario 0) was studied (Liu et al., 2020). Subsequently, a solution was proposed that is decentralized, the CA-UCB algorithm (Liu et al., 2021). CA-UCB finds a stable match in the APCK model by avoiding conflicts using the notion of a “plausible set.” The plausible set for a player $p_{i}$ at time-step $t$ is defined as $S_{i}(t):=\{a_{k}:p_{i}\succ_{a_{k}}p_{j},\text{ where }\bar{A}_{j}(t-1)=a_{k}\}$ i.e. the set of all arms $a_{k}$ such that at time-step $(t-1)$ another player $p_{j}\neq p_{i}$ successfully pulled the arm, but $p_{i}$ knows that $a_{k}$ prefers $p_{i}$ over $p_{j}$. This set also includes the arms the player successfully pulled (if any) and all the arms that were unmatched at the previous time-step. This makes it so that $p_{i}$ only attempts arms where there is a possibility of it successfully pulling the arm, avoiding conflicts. Then, $p_{i}$ attempts to pull the arm with the highest UCB value among $S_{i}(t)$ The UCB value tracked by $p_{i}$ for an arm $a_{k}$ is calculated as:
$\text{UCB}_{i}^{k}=\hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}$
where $\hat{\mu}_{i}^{k}$ is the empirical mean reward tracked by $p_{i}$ for $a_{k}$ up to time $t$, and $n_{i}^{k}(t)$ is the total number of samples that is used to calculate that empirical mean.
4. APKP Model and OCA-UCB
Note that the CA-UCB approach works only when both players and arms have full knowledge of the arms’ preferences, which players use to construct $S_{i}(t)$. When players lack this information, they must learn their positions in the arms’ preference orderings and recreate plausible sets. Recent game-theoretic approaches address this scenario. For instance, (Etesami and Srikant, 2024) allows proposers to maintain and update probability distributions over potential arms based on feedback from previous rounds. Accepted proposals increase the likelihood of future proposals to that arm, while rejections decrease it. These methods guarantee convergence to a stable matching without centralized coordination or full knowledge of preferences, achieving logarithmic regret in hierarchical (Etesami and Srikant, 2024) markets. Similarly, some approaches use mood-based state variables—like “content,” “discontent,” and “watchful”—to govern each player’s strategy at each time-step (Shah et al., 2024). Based on utility feedback, proposers decide whether to stick with their current match or experiment with new proposals. This ensures that over repeated interactions, proposers learn their preferences and the system converges to the proposer-optimal stable match, even in fully decentralized and information-limited markets. We present a much simpler approach to the solution which requires “collisions” with other players at most once for the players to learn arm preferences. We call it OCA-UCB (Optimistic CA-UCB) which exploits the fact that the arms still have full knowledge of their own preferences in order to achieve this.
Following the general structure of the CA-UCB algorithm, all players initially set their UCB estimates for the arms to $\infty$. However, now in addition to the UCB estimates, each player will also keep track of their position in the arms’ preference orderings. This can be done simply by having each player maintain two sets: $O^{k_{h}}_{(i)}$ the set of all players that $p_{i}$ believes are ranked higher than itself in $a_{k}$’s preference ordering, and $O^{k_{l}}_{(i)}$ which is the set of all players that $p_{i}$ believes are ranked lower than itself. Initially, for each player $p_{i}$, these sets are initialized to $O^{k_{h}}_{(i)}=\emptyset$ and $O^{k_{l}}_{(i)}=\left\{p_{j}\right\}_{j=1}^{N}\setminus p_{i}$. Thus, each player starts with the optimistic belief that it is the most preferred by each of the arms.
Notice, when updating plausible sets, the only information that $p_{i}$ requires is its’ relative positioning in $a_{k}$’s preference with respect to $p_{j}$ if $\bar{A}_{j}(t-1)=a_{k}$. This can be accomplished using the two sets described above. Since each player starts out with an optimistic view of itself in the arms’ preference orderings, conflicts are almost inevitable. This is because according to each player, all the arms are in the plausible set, even when they ought not to be. Suppose conflict occurs between two players, say $p_{i}$ and $p_{j}$, for an arm $a_{k}$ at time-step $t$, and player $p_{j}$ wins the conflict, then $p_{i}$’s belief with respect to $p_{j}$ and $a_{k}$ gets updated. More precisely, $O^{k_{h}}_{(i)}=O^{k_{h}}_{(i)}\cup\{p_{j}\}$ and $O^{k_{l}}_{(i)}=O^{k_{l}}_{(i)}\setminus\{p_{j}\}$. At future time steps, $p_{i}$ can use this information when considering $a_{k}$ in its plausible set. Functionally the rest of the algorithm is the same as CA-UCB. The only difference is that, at the end of each time-step, the players make use of the matching information and update $O^{k_{h}}_{(i)}$ and $O^{k_{l}}_{(i)}$ accordingly. The full algorithm is shown in Algorithm 1.
Algorithm 1 Optimistic Conflict Avoiding - UCB (OCA-UCB)
1:Input: $\lambda$, $H$ 2:Output: A stable match 3:for $t\in\{1,\cdots,H\}$ do 4:     for $i\in\{1,\cdots,N\}$ do 5:         if $t==1$ then 6:              $A_{i}(t)\leftarrow a_{k},k\sim\text{Uniform}(\left[K\right])$ 7:         else 8:              $D_{i}(t)\sim Ber(\lambda)$ 9:              if $D_{i}(t)==0$ then 10:                  $A_{i}(t)\leftarrow$ GetBestArm($p_{i},t$) 11:              else 12:                  $A_{i}(t)\leftarrow A_{i}(t-1)$ 13:              end if 14:         end if 15:     end for 16:     for $k\in\{1,\cdots,K\}$ do 17:         $p_{w}\leftarrow$ ResolveRequests($a_{k},t$) 18:         if $p_{w}\neq\emptyset$ then 19:              $r_{p_{w}},r_{a_{k}}\leftarrow$ Sample-Rewards($p_{w},a_{k}$) 20:              Arm-Update-Belief($a_{k},p_{w},r_{a_{k}}$) 21:              Player-Update-Belief($p_{w},a_{k},r_{p_{w}}$) 22:              $L\leftarrow\left\{p_{j}:A_{j}(t)=a_{k}\right\}\setminus\left\{p_{w}\right\}$ 23:              for $p_{l}\in L$ do 24:                  UpdateConflictBelief($p_{l},p_{w},a_{k}$) 25:              end for 26:         end if 27:     end for 28:end for 1:procedure GetBestArm ($p_{i}$, $t$) 2:     $S_{i}(t)\leftarrow\emptyset$ 3:     for $k\in\{1,\cdots,K\}$ do 4:         $p_{\bar{p}}\leftarrow$ $p_{j}\text{ s.t. }\bar{A}_{j}(t-1)=a_{k}$ 5:         if $p_{\bar{p}}\in O_{(i)}^{k_{l}}\vee p_{\bar{p}}=\emptyset$ then 6:              $S_{i}(t)\leftarrow S_{i}(t)\cup\{a_{k}\}$ 7:         end if 8:     end for 9:     $k\leftarrow\underset{a_{j}\in S_{i}(t)}{\mathop{\mathrm{arg\,max}}\limits}\left[\text{UCB}_{i}^{j}\right]$ 10:     return $a_{k}$ 11:end procedure 1:procedure UpdateConflictBelief($p_{i},p_{j},a_{k}$) 2:     $O_{(i)}^{k_{l}}\leftarrow O_{(i)}^{k_{l}}\setminus p_{j}$ 3:     $O_{(i)}^{k_{h}}\leftarrow O_{(i)}^{k_{h}}\cup\{p_{j}\}$ 4$\triangleright$ $p_{i}$ lost conflict with $p_{j}$ for $a_{k}$ 5:end procedure
4.1. Proof of OCA-UCB Convergence
As Algorithm 1 progresses, each of the players will have complete sets that represent the arms’ true preferences with respect to the other players. In fact, we show theoretically that this approach to the solution does not affect the convergence guarantee of the CA-UCB algorithm in the APCK Model in Theorem 4.1.
{restatable}
theoremSimpleTheorem Under the belief update scheme described above, OCA-UCB shares CA-UCB’s guarantee of convergence to stability.
First, let us define some terminology. A triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ is inconsistent if $p_{i}$ believes $p_{i}\succ_{a_{k}}p_{j}$ but the opposite is true. Then, let us define a plausible set constructed by $p_{i}$ to be inconsistent with respect to $p_{j}$ and $a_{k}$ if $a_{k}\in S_{i}(t)$, denoted by $\hat{S}_{ij}^{k}(t)$. Finally, let us define an inconsistent plausible set as being inconsequential at time $t$ if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{i}^{m}:a_{m}\in S_{i}(t)\}\neq k$ i.e. $p_{i}$ does not attempt $a_{k}$ OR if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{j}^{m}:a_{m}\in S_{j}(t)\}\neq k$ i.e. $p_{j}$ does not attempt $a_{k}$. We begin the proof of the convergence guarantee by first presenting two lemmas:
Lemma 4.0.
At each time-step, inconsistent sets $\hat{S}_{ij}^{k}(t)$ are either inconsequential or get resolved such that $S_{i}(t+1)$ is no longer inconsistent.
Proof.
Consider an inconsistent triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ such that $\bar{A}_{j}(t-1)=a_{k}$. Then, at time $t$, $\hat{S}_{ij}^{k}(t)$- the plausible set constructed by $p_{i}$ will be inconsistent. One of three things must happen at time $t$ with regard to $\hat{S}_{ij}^{k}(t)$. (1) $p_{j}$ attempts $a_{k}$ either due to $\lambda$, or because $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{j}^{m}:a_{m}\in S_{j}(t)\}=k$. $p_{i}$ also attempts $a_{k}$ because $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{i}^{m}:a_{m}\in\hat{S}_{ij}^{k}(t)\}=k$. This leads to a conflict, resulting in the inconsistency getting resolved via $p_{i}$ receiving feedback about matching information. (2) $p_{j}$ attempts $a_{k}$ but $p_{i}$ does not. This implies $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{i}^{m}:a_{m}\in\hat{S}_{ij}^{k}(t)\}\neq k$, which in turn implies $\hat{S}_{ij}^{k}(t)$ is inconsequential with respect to $p_{j}$ and $a_{k}$. (3) $p_{j}$ does not attempt $a_{k}$ but $p_{i}$ does. This implies $\hat{S}_{ij}^{k}(t)$ is inconsequential with respect to $p_{j}$ and $a_{k}$. However, two further distinct sub-cases follow from this. First, if $\bar{A}_{i}(t)=a_{k}$ then $p_{i}$ was the most preferred among $a_{k}$’s incoming pull requests. The plausible set constructed at $(t+1)$ is no longer inconsistent with respect to $p_{j}$ and $a_{k}$ because $a_{k}$ will be included in it by virtue of it being successfully pulled by $p_{i}$ at $t$, and not because of $p_{i}$’s belief $Q_{ij}^{k}$. And second, if $\bar{A}_{i}(t)\neq a_{k}$ then $a_{k}$ will not be included in plausible set constructed at $(t+1)$ as $p_{i}$ will lose a conflict with $M(a_{k})$ and $p_{i}$’s belief will be updated with respect to $M(a_{k})$ and $a_{k}$ if it has not already. ∎
Lemma 4.0.
Eventually all inconsistent plausible sets constructed are resolved or are inconsequential.
Proof.
First, begin by noting that there are only finitely many inconsistent triplets in any given APKP model. Then, a simple recursive argument will suffice to show that each one gets resolved or becomes inconsequential. The first and second cases from Lemma 4.1 are base cases. The first case corrects $p_{i}$ belief altogether, and in the second case the incorrect belief never gets used in any significant way. In the first sub-case of case three, if $a_{k}$ is in-fact $p_{i}$’s best possible achievable arm, then it will continue getting matched with $a_{k}$ regardless of it’s wrong belief (inconsequential w.r.t $a_{k}$). Nonetheless, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ exists, then we can apply Lemma 4.1 recursively. And finally, in the second sub-case of case three from Lemma 4.1, $p_{i}$ loses the conflict to some other player which implies that some belief gets updated resulting in the resolution of some triplet. However, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{i,j}^{k}=(p_{i},p_{j},a_{k})$ exists, then again we can apply Lemma 4.1 recursively. ∎
The proof of Theorem 4.1 is the immediate consequence of lemmas 4.1, and 4.2. Once all the inconsistent plausible sets are either resolved, or are inconsequential, the best arm picked from each player’s plausible set will be the same as the one picked from CA-UCB’s plausible set. This implies that this approach to updaing beliefs has no effect on CA-UCB algorithm’s guarantee on convergence. ∎
We present empirical evidence on the performance of OCA-UCB in Section 8.1.
5. APU Model and PCA-SCA
The previous section looked at the APKP model where the agents did not have access to arms’ preferences. Instead, the players relied on the fact that the arms had perfect knowledge about their own preferences and used deterministic feedback obtained to recreate something similar to the plausible set from the CA-UCB algorithm. However, in the APU model, even arms lack knowledge of their preferences, making feedback received from them noisy, invalidating the approach from the previous section.
Given the issue of unreliable feedback associated with conflict results, we must formulate an approach in which agents can pick the best possible arms while avoiding conflicts. To this end, we use the notion of highest expected reward to determine agent proposals. Each player tracks its estimated probability of winning a conflict against other players. Players choose the arm that maximizes the product of reward estimates with conflict win probability estimates.
Concurrently, in our algorithm, each arm maintains an empirical reward estimate for each player along with a confidence interval. These confidence intervals serve two purposes: first, when the confidence intervals for a given set of players are disjoint, the arm can deterministically identify the player with the highest expected reward. Second, when the intervals overlap, the arm resolves the conflict by selecting a player uniformly at random among those with overlapping intervals. This allows us to maintain the convergence guarantees of our approach. We detail the algorithmic implementation of this approach in Algorithm 2.
Algorithm 2 Probabilistic Conflict Avoiding - Simultaneous Choice Algorithm (PCA-SCA)
1:Input: $\lambda$, $\kappa$, $H$ 2:Output: A stable match 3:for $t\in\{1,\cdots,H\}$ do 4:     for $i\in\{1,\cdots,N\}$ do 5:         if $t==1$ then 6:              $A_{i}(t)\leftarrow a_{k},k\sim\text{Uniform}(\left[K\right])$ 7:         else 8:              $D_{i}(t)\sim Ber(\lambda)$ 9:              if $D_{i}(t)==0$ then 10:                  $A_{i}(t)\leftarrow$ GetBestArm($p_{i},t$) 11:              else 12:                  $A_{i}(t)\leftarrow A_{i}(t-1)$ 13:              end if 14:         end if 15:     end for 16:     for $k\in\{1,\cdots,K\}$ do 17:         $p_{w}\leftarrow$ ResolveRequests($a_{k},t$) 18:         if $p_{w}\neq\emptyset$ then 19:              $r_{p_{w}},r_{a_{k}}\leftarrow$ Sample-Rewards($p_{w},a_{k}$) 20:              Arm-Update-Belief($a_{k},p_{w},r_{a_{k}}$) 21:              Player-Update-Belief($p_{w},a_{k},r_{p_{w}}$) 22:              $L\leftarrow\left\{p_{j}:A_{j}(t)=a_{k}\right\}\setminus\left\{p_{w}\right\}$ 23:              for $p_{l}\in L$ do 24:                  Update-Win-Probability($p_{l},p_{w},a_{k}$) 25:              end for 26:         end if 27:     end for 28:end for 1:procedure GetBestArm ($p_{i}$, $t$) 2:    for $k\in\{1,\cdots,K\}$ do 3:        $p_{\bar{p}}\leftarrow$ $p_{j}\text{ s.t. }\bar{A}_{j}(t-1)=a_{k}$ 4:        REWARDS$[k]\leftarrow\mathcal{R}_{i}^{k}$triangleright Reward estimate 5:        $P[k]\leftarrow f\left(\hat{P}_{i\bar{p}}^{k}\right)$triangleright Optimism Function 6:    end for 7:    $j\leftarrow\arg\max\left(\text{REWARDS}\circ\text{P}\right)$ 8:    return $a_{j}$ 9:end procedure 1:procedure ResolvePullRequests($a_{k},t$) 2:    $A_{rq}\leftarrow\left\{p_{i}:A_{i}(t)=a_{k}\right\}$$\triangleright$ pull requests 3:    $O_{inf}\leftarrow\emptyset$triangleright players with $\infty$ reward estimate 4:    $O_{bst}\leftarrow\emptyset$triangleright players with best overlapping reward estimate 5:    $I_{L},I_{U}\leftarrow\infty,-\infty$triangleright Keep track of best reward interval 6:    for $p_{i}\in A_{rq}$ do 7:        $\mathcal{R}_{i}^{U}\leftarrow{\tiny\text{Upper bound on reward estimate for }p_{i}}$ 8:        $\mathcal{R}_{i}^{L}\leftarrow{\tiny\text{Lower bound on reward estimate for }p_{i}}$ 9:        if $\mathcal{R}_{i}^{U}=\infty$ then 10:           $O_{inf}=O_{inf}\cup\{p_{i}\}$ 11:           Continue 12:        end if 13:        if $\mathcal{R}_{i}^{L}>I_{U}$ then $\triangleright$ Better interval 14:           $I_{H}\leftarrow\mathcal{R}_{i}^{U}$, $I_{L}\leftarrow\mathcal{R}_{i}^{L}$ 15:           $O_{bst}=\{p_{i}\}$ 16:        else$(I_{U}\leq\mathcal{R}_{i}^{L})\wedge(\mathcal{R}_{i}^{U}\leq I_{L})$triangleright Overlap 17:           $O_{bst}=O_{bst}\cup\{p_{i}\}$ 18:        end if 19:    end for 20:    if $O_{inf}\neq\emptyset$ then 21:        return $\sim\text{Uniform}(O_{inf})$ 22:    else 23:        return $\sim\text{Uniform}(O_{bst})$ 24:    end if 25:end procedure
The algorithm is parameterized by $\lambda\in[0,1)$, and $\kappa\geq 1$. The first parameter $\lambda$ is used to introduce a random delay mechanism to reduce the likelihood of conflicts (Liu et al., 2021; Kong et al., 2022). This is controlled in line 8 of PCA-SCA where each player draws a Bernoulli Random Variable $D_{i}(t)$ with expectation $\lambda$. If $D_{i}(t)=0$, the player attempts the arm with the highest expected reward, else it will attempt the arm that it did in the previous time step. The second parameter $\kappa$ is used to control the level of optimism that the player uses when determining the probability of it winning a conflict against another player for an arm. This is used in line 5 of procedure GetBestArm. We will shortly detail the specifics of how this level of optimism is determined. Initially all reward estimates are set to infinity, and all win conflict probabilities are set to 1. First let us go over how the arms pick the players at each round.
Arms’ Choice of Players:
At each time step, players submit proposals to the arms according to a specified protocol (detailed next). Each arm, upon receiving one or more proposals, must select exactly one proposing player. Because the arms initially do not know their preferences, each arm maintains beliefs about the payoff of each player and confidence intervals for those beliefs.
Each arm processes the incoming proposals as follows. If at least one player’s reward estimate is still infinite (i.e., uninitialized), the arm selects uniformly at random from those players (Line 21 of ResolvePullRequests). Otherwise, the arm identifies the player with the highest upper confidence bound and checks whether other players’ confidence intervals overlap with this best bound (Line 16). If multiple intervals overlap, the arm selects uniformly at random from among those candidates (Line 23); if only one player remains with the highest bound, it selects that player.
Players’ Choice of Arms:
As in CA-UCB, each player begins with all arm reward estimates set to $\infty$. However, players also maintain probability estimates $\hat{P}_{ij}^{k}$, which represent player $p_{i}$’s belief of winning a conflict against player $p_{j}$ for arm $a_{k}$. These estimates are updated based on the ratio of conflicts won to total conflicts encountered and are initialized to 1.
Over time, players gather conflict outcomes and adjust $\hat{P}_{ij}^{k}$ accordingly. At each time step, $p_{i}$ selects the arm that maximizes its expected reward:
$\arg\max_{k}\Bigl{[}\hat{P}_{ij}^{k}\cdot\mathcal{R}_{i}^{k}\Bigr{]},$
where $j$ is the index of the player who pulled arm $k$ in the previous step, and $\mathcal{R}_{i}^{k}$ is $p_{i}$’s current reward estimate for arm $a_{k}$.
Remember that while the arms’ confidence intervals overlap, they select players uniformly at random, rendering conflict feedback noisy. In expectation, $\hat{P}_{ij}^{k}$ converges to $0.5$ during this phase. Once the intervals become disjoint, selection becomes deterministic, allowing $\hat{P}_{ij}^{k}$ to approach its true value. However, a player with an “unlucky” estimate that drifts below $0.5$ may prematurely abandon that arm. To encourage continued exploration—even under unfavorable conflict estimates—we introduce the optimism function:
(1) $f(x)=\begin{cases}\displaystyle\frac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5,\end{cases}$
where $\kappa\geq 1$ is a hyper-parameter that controls the level of optimism in the algorithm. Equation 1 is employed to modulate optimism due to its favorable behavior with respect to each $\hat{P}_{ij}^{k}$. Specifically, if $\hat{P}_{ij}^{k}>0.5$, the player is fully optimistic about its win chances; however, when $\hat{P}_{ij}^{k}$ falls below 0.5, the function $f(\hat{P}_{ij}^{k})$ boosts the estimate to 1 and then decreases it gradually, with the rate of decrease governed by $\kappa$. This behavior is illustrated in Figure 6 (in the Appendix), which plots $f(x)$ for different values of $\kappa$.
As $\kappa\to\infty$, $f(\hat{P}_{ij}^{k})$ approaches 1 regardless of the value of $\hat{P}_{ij}^{k}$, whereas $\kappa=1$ yields an almost linear, yet optimistic, dropoff. As players begin receiving deterministic feedback, $\hat{P}_{ij}^{k}$ converges to the true win probability; if that value is 1, then any estimate with $\hat{P}_{ij}^{k}\geq 0.5$ is boosted to 1 by the function, while if the true probability is 0, the function decreases the estimate slowly—again, at a rate determined by $\kappa$. Thus, the choice of $\kappa$ should reflect how aggressively one wishes to maintain optimism.
With this information, at each subsequent time step, the players will attempt to pull the arm that maximizes the product of the reward estimate and the optimistic estimate of the probability of winning (line 7 of procedure GetBestArm).
Then, at the end of each time step, the arms will pick the player that it believes will give the highest reward (similar to how arms behave in the simultaneous choice mechanism of (Das and Kamenica, 2005)) – with the added caveats mentioned above. A player and arm both receive a reward and update their beliefs when an arm is pulled successfully. No reward is given if an arm is not pulled or if a pull request is not received. Players update their beliefs about winning after all matches are made.
Finally, the structure of our algorithm accommodates multiple methods for updating beliefs about expected rewards and conflict win probabilities. In this work, we consider two approaches: UCB, as employed in (Liu et al., 2021), and a Thompson Sampling variant, as presented in (Kong et al., 2022). We prove that using UCB for belief maintenance ensures convergence to stability and support this claim with simulation results. We present simulation results for the Thompson Sampling variant, leaving a theoretical convergence proof for future work.
6. Using UCB in PCA-SCA (PCA-UCB)
In this section, we look at using the Upper Confidence Bound (UCB) approach to keep track of reward estimates, as used in the multi-armed bandit literature by (Auer et al., 2002). We use UCB in the PCA-SCA algorithm and refer to it as PCA-UCB. Players must keep track of beliefs about payoffs and the probability of winning conflicts. Equation 2 describes how beliefs about reward payoffs are estimated by the players.
(2) $\text{UCB}_{i}^{k}(t)=\begin{cases}\infty&\qquad\text{ if }n_{i}^{k}(t)=0\\ \hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t-1)}}&\qquad\text{ otherwise}\end{cases}$
where $n_{i}^{k}(t)$ is the number of times player $p_{i}$ has pulled arm $a_{k}$ at time-step $t$, and $\hat{\mu}_{i}^{k}$ is the empirical mean tracked by player $p_{i}$ for arm $a_{k}$. Similarly, the arms keep track of the confidence intervals for the rewards offered by the players. This is done by having the arms keep track of both the UCB and the Lower Confidence Bound (LCB) associated with each player:
(3) $\begin{array}[]{rcl}\text{LCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)-c_{i}^{k}(t),\\ \text{UCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)+c_{i}^{k}(t)\end{array}$
where $c_{i}^{k}(t)=\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}$. Similar to the players, the LCB and the UCB for the arms is initialized to $\infty$- ∞ and $\infty$ respectively. The other quantity that the players keep track of is the probability of conflict wins:
(4) $\hat{P}_{ij}^{k}(t)=\begin{cases}1&\qquad\text{ if }n_{ij}^{k}=0\\ \frac{w_{ij}^{k}(t)}{N_{c_{ij}}^{k}(t)}&\qquad\text{ otherwise}\end{cases}$
where $\hat{P}_{ij}^{k}(t)$ represents player $p_{i}$’s belief about the probability of winning a conflict against player $p_{j}$ for arm $a_{k}$ at time $t$. $w_{ij}^{k}$ is the number of times player $p_{i}$ has won this conflict and $N_{c_{ij}}^{k}(t)$ is the total number of times this conflict has happened thus far. At the end of each time step, belief estimates are updated once rewards are sampled and matching information is made public. If a player wins a conflict, the corresponding probability estimate increases, otherwise, it decreases. This is handled by the Update-Win-Probability function in Algorithm 2.
Before presenting the sketch of the proof of convergence, we will first provide some preliminary definitions and notation. First, in our analysis, we assume without loss of generality that players are ordered so that $\mu_{i}^{k}>\mu_{j}^{k}$ when $p_{i}\succ_{a_{k}}p_{j}$. We say that the confidence intervals for arm $a_{k}$ are disjoint if, for every pair $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$,
$\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t).$
Conversely, if for some $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$ we have
$\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t),$
then the confidence intervals for those two players are said to be overlapping. As mentioned before, when an arm’s confidence intervals overlap for a set of players, the conflict feedback is provided uniformly at random. Specifically, the arm selects the proposal with the highest UCB estimate; if there is any overlap with another player’s interval, it chooses uniformly at random from the contenders. Since pairwise win probability estimates are only updated when one of the two members of the pair actually wins the arm at that time (thus conflicts between more than two players do not confound the estimate), the win probability estimates for losing players concentrate around $0.5$. Once the intervals become disjoint, however, the feedback becomes deterministic. Let us now define this deterministic feedback event:
Definition 0 (Deterministic Feedback Event $\alpha(t)$).
For any time $t$, define
$\alpha(t)=\Bigl{\{}\forall\,k\in[K],\,\forall\,i,j\text{ with }\mu_{i}^{k}>\mu_{j}^{k}:\;\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t)\Bigr{\}}.$
Thus, $\alpha(t)$ is the event that, for every arm, the players’ confidence intervals are disjoint at time $t$. When $\alpha(t)$ holds, the arm’s feedback is deterministic.
In CA-UCB, players had direct access to arms’ preferences, whereas in OCA-UCB, the arms’ preferences were learned deterministically. In the context of the APKP model, one may view an arm $a_{k}\in S_{i}(t)$ as having a win probability of 1 (with all other arms having a win probability of 0), thereby considering arms with a win probability of 1 as “plausible”. In the APU model, however, plausibility must be interpreted more generally as any arm with a non-zero probability of being won.
Definition 0 (Inconsistent Triplet).
We now define an inconsistent triplet in the APU model. Let $P_{ij}^{k}$ denote the true probability of player $p_{i}$ winning against player $p_{j}$ in a conflict for arm $a_{k}$, and let $\hat{P}_{ij}^{k}$ be the corresponding empirical probability. The triplet $(p_{i},p_{j},a_{k})$ is said to be inconsistent if
[$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
i.e., the player’s empirical belief about its win probability in the conflict deviates significantly (by at least $\epsilon$) from the true outcome—either overestimating a win or underestimating a loss.
Next, we define a inconsequential triplet
Definition 0 (Inconsequential Triplet).
An inconsistent triplet is considered inconsequential if there exists an alternative arm that offers a better reward prospect. Formally, consider two triplets $(p_{i},p_{j},a_{k})$ and $(p_{i},p_{j}^{\prime},a_{k}^{\prime})$. If, at time $t$, the triplet $(p_{i},p_{j},a_{k})$ is inconsistent but $\text{UCB}_{i}^{k^{\prime}}>\text{UCB}_{i}^{k}$ and the triplet $(p_{i},p_{j}^{\prime},a_{k}^{\prime})$ is consistent (i.e., arm $a_{k^{\prime}}$ is a better prospect for $p_{i}$), then player $p_{i}$ will select $a_{k^{\prime}}$. In this case, the inconsistency in $(p_{i},p_{j},a_{k})$ is rendered inconsequential.
A triplet is considered consequential if it is not inconsequential. And finally, for any pair of players $p_{i}$, $p_{j}$, and $a_{k}$, define:
$\begin{array}[]{rcl}\Delta_{ij}^{k}&=&\mu_{i}^{k}-\mu_{j}^{k}\quad\text{for }\mu_{i}^{k}>\mu_{j}^{k},\\ \Delta_{\min}&=&\underset{i\neq j,\forall k}{\min}\lvert\mu_{i}^{k}-\mu_{j}^{k}\rvert,\\ n_{ij}^{k}(t)&=&\min\{n_{i}^{k}(t),n_{j}^{k}(t)\}.\end{array}$
Note that, without loss of generality, in our models, we assume $\Delta_{\min}=1$. With these preliminaries out of the way, we will now proceed to provide a sketch of the proof that proves that using UCB in the PCA-SCA algorithm yields convergence to market stability. A full proof can be found in Appendix A.
6.1. Sketch of the Proof of Convergence to Stability Using PCA-UCB
Consider an APU model with $K$ arms and $N$ players. Our approach is to show that (i) each arm learns its true preference ordering over players in finite time, (ii) once the arms’ estimates are sufficiently accurate, the conflict feedback becomes deterministic so that the players’ empirical win probability estimates (that is, the estimated probability of winning a conflict against a specific other player for a specific arm) converge to their true values (either 0 or 1), and (iii) as a consequence, the execution of our PCA-UCB algorithm becomes identical to that of CA-UCB in the long run.
The first step is to establish that each arm maintains an empirical reward estimate for every player together with a confidence interval that shrinks as the number of samples increases. By bounding the probability that the confidence intervals of two players overlap—that is, that the lower confidence bound of a higher-mean player exceeds the upper confidence bound of a lower-mean player—we show that the probability of any overlap decays exponentially with the number of observations. We denote the event that all arms have learned their preferences by time $t$ as $\alpha(t)$.
Conditioned on the event $\alpha(t)$, the feedback that players receive during conflicts becomes deterministic. In the early stages, when the arms’ confidence intervals may overlap, the feedback is random and the players’ win estimates concentrate around $0.5$. To prevent premature abandonment of an arm due to these fluctuations, our algorithm incorporates an optimism function that boosts the win probability estimates. Once the feedback is deterministic, each conflict yields a noise-free observation, and a straightforward averaging argument shows that the error in the players’ empirical win estimates decreases proportionally to $\frac{T}{T+t}$ after time $T$.
Finally, by combining the bounds on the probability that the arms’ confidence intervals overlap with the convergence rate of the win estimates, we conclude that the overall failure probability vanishes over time. In this asymptotic regime, the decision-making process of PCA-UCB exactly replicates that of CA-UCB. Now, we will provide the sketch of the proof of each of the theorems.
{restatable}
theoremTheoremOne[Finite-Time Learning of Arms’ Preferences] In an instance of the APU model with $K$ arms and $N$ players running PCA-UCB, at time $t$, with probability at least $1-\delta(t)$, for every arm, the confidence intervals that the arm tracks for each player are disjoint. Here $\delta(t)$ is :
$\delta(t)=K\cdot\frac{N^{2}}{2}\cdot\exp\left(-\frac{n(t)\left(\Delta_{\text{min}}-2c(t)\right)^{2}}{4}\right)$
where $n(t)=\underset{\begin{subarray}{c}i\in N\\ k\in K\end{subarray}}{\min}\left\{n_{i}^{k}(t)\right\}$, $\Delta_{\min}=\underset{\begin{subarray}{c}i,j\in N\\ k\in K\end{subarray}}{\min}\left\{\Delta_{ij}^{k}\right\}$, and $c(t)=\sqrt{\frac{3\ln(t)}{2n(t)}}$.
To prove Theorem 6.1, we first analyze the overlap probability for a single pair of players on a fixed arm. By writing the event of overlapping confidence intervals as
$\hat{\mu}_{i}(t)-\hat{\mu}_{j}(t)\leq c_{i}(t)+c_{j}(t),$
and by applying a concentration inequality (specifically, a Hoeffding bound for sub-Gaussian random variables), we obtain an exponentially decaying bound on the probability that the intervals for any pair do not separate. A union bound over all player pairs and over all $K$ arms then shows that the probability that any arm has overlapping intervals is bounded by $\delta(t)$. In other words, as the number of samples $n(t)$ grows, the event $\alpha(t)$—where every arm’s confidence intervals are disjoint—occurs with high probability.
{restatable}
theoremTheoremTwo[Player Win Estimate Convergence] Assume that event $\alpha$ has occurred by time $T$. Then, at time $T+t$, with probability at least $1-\tau(T)$ for all consequential conflicts $(p_{i},p_{j},a_{k})$ (defined above):
$\lvert\hat{P}_{ij}^{k}(T+t)-p_{ij}^{k}\rvert\leq\frac{T}{T+t}$
Here $\tau(t)$ is:
$\frac{KN^{2}}{2}\exp\left(-2\,N_{c}(T)\,\left(\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}\right)^{2}\right)$
where $N_{c}(t)$ is the minimum number of conflicts across all triplets.
Theorem 6.1 is proved by first noting that prior to $\alpha$ the feedback provided by the arms is random, leading to win estimates that concentrate around $0.5$. To ensure that players continue to sample even when their estimates are slightly off, we introduce an optimism function $f(x)$ that amplifies estimates near $0.5$. Once $\alpha$ has occurred, the conflict outcomes become deterministic so that each conflict provides a true (noise-free) observation. Using a simple averaging argument, one can then show that the error in the win estimate decays as $\frac{T}{T+t}$. Moreover, any triplet for which the win estimate does not converge sufficiently fast is rendered inconsequential to the player’s decision-making. A careful matching of the tail probabilities of the concentration inequalities yields the bound $\tau(t)$.
{restatable}
theoremTheoremThree[Identical Execution to CA-UCB] After time $T+t$, the execution of PCA-UCB is identical to that of CA-UCB.
The proof of Theorem 6.1 follows directly from Theorems 6.1 and 6.1. Since the probability of the arms’ confidence intervals overlapping (and thus producing non-deterministic feedback) decays to zero, and because the players’ win estimates converge to their true values with high probability, the overall probability that any “failure” occurs (i.e., either an arm has overlapping intervals or a consequential triplet’s win estimate is inaccurate) tends to zero as time increases. Consequently, the decision rules used by PCA-UCB reduce to those of CA-UCB in the limit.
In summary, our proof strategy first ensures that each arm learns its preference over players via disjoint confidence intervals (Theorem 6.1). This in turn guarantees that the conflict feedback becomes deterministic, allowing the players’ win estimates to converge (Theorem 6.1). Finally, these results imply that the algorithm’s action selection mirrors that of CA-UCB, yielding asymptotic equivalence (Theorem 6.1).
Note that convergence in practice occurs much faster than would be implied by the worst-case analysis. While we derive a bound based around the event $\alpha(T)$, the design of the algorithm ensures that once the confidence intervals for an associated triplet become disjoint, the players immediately receive deterministic feedback—well before the occurrence of $\alpha(T)$. The sooner the players obtain deterministic conflict feedback, the faster the algorithm converges to stability. Moreover, players do not have to wait for $\alpha(T)$ or for the conflict probabilities to converge before acquiring reward estimates for the arms. Consequently, even though the algorithm ultimately replicates CA-UCB, the players begin with some prior information regarding the arms’ rewards, implying that convergence in practice could be significantly tighter, which is consistent with our experimental findings below.
7. Using Thompson Sampling in PCA-SCA (PCA-TS)
Thompson Sampling (Thompson, 1933) (TS) is another approach to solving the MAB problem which has seen a recent resurgence in the literature (Agrawal and Goyal, 2012; Chapelle and Li, 2011; Kong et al., 2022). In the PCA-SCA algorithm, TS can be used to keep track of both the beliefs about rewards by both players and arms. Henceforth, we will refer to this approach as the PCA-TS algorithm. The players and arms keep track of the same information as in UCB, i.e. the total reward obtained for an arm/player, the total number of pulls of an arm, and the total number of conflict wins against each player. The distinction is in how this information is used to update beliefs.
First, players use the same estimates for conflict wins as shown in Equation 4, and subsequently use Equation 1 for an optimistic estimate of winning a conflict. As for the rewards, we assume the variance ($\sigma^{2}=1$) in rewards of the arms and the players are known. Then, we update the mean and precision ($\tau=\frac{1}{\sigma^{2}}$) as:
(5) $\mu_{\text{new}},\tau_{\text{new}}=\frac{\tau_{0}\mu_{0}+\tau\sum_{i=1}^{n}x_{i}}{\tau_{0}+n\tau},\tau_{0}+n\tau$
where $\{\mu_{\text{new}},\tau_{\text{new}}\}$ are the new mean and precision, $\{\mu_{0},\tau_{0}\}$ are the old mean and precision, $\tau=1$ is the known true precision, and $\sum x_{i}$ is the sum of rewards for the agent in question. We can use appropriate lower and upper bounds on the confidence intervals, $\mathcal{R}_{i}^{L}(t),\mathcal{R}_{i}^{U}(t)$.111 In our experiments below, we use $\mathcal{R}_{i}^{L}(t)=\mu_{0}-\frac{1}{\tau_{0}},\mathcal{R}_{i}^{U}(t)=\mu_{0}+\frac{1}{\tau_{0}}$ where $\frac{1}{\tau_{0}}$ represents the standard deviation in belief. These are the lower and upper confidence bounds respectively (Lines 7, 8 respectively in ResolvePullRequests of Algorithm 2). When an agent needs an estimate for a reward for a particular player or an arm, it samples from $\mathcal{N}(\mu_{0},\frac{1}{\tau_{0}})$ where $\mu_{0}$ and $\tau_{0}$ is the corresponding mean and precision tracked by the agent for the specific player/arm. We provide simulation results of using PCA-TS for different market sizes and varied preference heterogeneity in Section 8.
8. Simulation Results
We run simulation experiments where (1) preferences are uniformly random on both sides of the market with varying market sizes $N=K\in\{5,10,15,20\}$; (2) player preference heterogeneity is varied with market size $N=K=10$. The maximum reward an agent can get is $(K+1)$, when matched with its most preferred arm, and the minimum is 1, when matched with the least preferred. The reward gaps (of the means) between consecutively ranked agents are kept constant at $\Delta=1$. To control the level of heterogeneity in player preferences, we use a method due to (Ashlagi et al., 2017), as also used by (Liu et al., 2021). The level of heterogeneity is controlled by a parameter $\beta\in[0,10,100,1000]$. Heterogeneity in preferences decreases with $\beta$, and as $\beta\to\infty$ all players have the same preferences with probability 1. To translate $\beta$ to preference orderings we sample the mean reward $\mu_{i}^{k}$ of arm $a_{k}$ for player $p_{i}$ using the following process:
(7) $\begin{split}x_{k}&\stackrel{{\scriptstyle i.i.d}}{{\sim}}\text{Uniform([0,1])}\\ \epsilon_{k,i}&\stackrel{{\scriptstyle i.i.d}}{{\sim}}\text{Logistic(0, 1)}\\ \overline{\mu}_{i}^{k}&=\beta\cdot x_{k}+\epsilon_{k,i}\\ \mu^{k}_{i}&=\#\{j:\overline{\mu}^{j}_{i}\leq\overline{\mu}_{i}^{k}\}\end{split}$
These random utilities $\mu_{i}^{k}$ are mapped so that the gap in rewards between consecutively ranked arms is kept constant at $\Delta=1$. The hyper-parameter $\lambda$ is set to $0.9$ and $\kappa$ is set to 10 (in the context of PCA-SCA). [^2] There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets (Ackermann et al., 2008). The method we use corresponds to uncoordinated markets when $\beta$ is low. However, coordinated markets are different, with correlation defined on the edges of the matching graph rather than the vertices. We find that the convergence time in these edge-correlated cases is actually comparable to when preferences are uniformly random, compared with node-correlation, where convergence time increases (see Appendix).
To objectively quantify the quality and rate of convergence, we introduce a notion of a “convergence proxy”: It measures, in sliding windows of size $\mathcal{X}$, the percentage of time-steps where market stability was greater than $\theta$. It effectively measures the tendency of a market to converge within the next $\mathcal{X}$ time-steps. We calculate the “convergence proxy” line for PCA-UCB vs PCA-TS for random preferences and varied market sizes. This allows us to objectively visualize the convergence behavior of the two approaches and draw conclusions about their behaviors.
We run simulations in each of the aforementioned scenarios 100 times and monitor the market state. More specifically, we take a snap-shot of the maximum player regret and market stability every 10 time-steps for each experiment and average the results over all the runs. To evaluate the trend in the quantities we study, Loess smoothing was run on the data points to yield the graphs presented in the following sections.
8.1. OCA-UCB in the APKP model
Recall that, in this scenario, the players must learn their own position in the arms’ preference order and their own preferences. We compare the performance of this model with the APCK model using CA-UCB. Both models were run with their respective algorithms, with 6000 time steps for random preferences and 3000 time steps for varied preference heterogeneity. The results are shown in Figure 1.
Figure 1. The left sub-figures show the APCK models using CA-UCB and the right sub-figures show the APKP model using OCA-UCB. Results are averaged over 100 runs. The first row varies market size with uniformly random preferences, the second row varies player preference heterogeneity while keeping market size constant at $N=K=10$. OCA-UCB converges to stability even with uncertainty on both sides, though it takes longer (expected because of increased complexity)
The figures in the first row of Figure 1 show results from the experiments where agents’ preferences are uniformly random. We can see that OCA-UCB converges to stability in the APKP model, with increasing market sizes corresponding to slower convergence. This trend is essentially the same as the convergence of CA-UCB in the APCK model, albeit OCA-UCB convergence in the APKP model is slightly slower. Thus, the OCA-UCB algorithm converges to stability in the APKP model almost as fast as the CA-UCB algorithm does in a model where all arm preferences are common knowledge.
The second row in Figure 1 details experimental results for varying player preference heterogeneity. As can be seen from the APCK Model running CA-UCB, decrease in player preference heterogeneity in the APCK model does not have any effect on the convergence of CA-UCB. This trend carries over to OCA-UCB in the APKP model, with convergence slightly slower. This shows that in the setting with varying levels of player preference heterogeneity, OCA-UCB does not show any noticeable difference in the rate at which the markets converge to stability.
In combination with the theoretical results above, our results show that learning to converge to stability is not noticeably harder when arm preferences are not common knowledge, as long as they are known to the arms.
8.2. PCA-UCB and PCA-TS in the APU Model
We now turn to the considerably more challenging setting where both the players and the arms do not have access to any preference information. Recall that in our approach, where at each time-step the players attempt the arm with the highest expected reward using the optimism function. As the arms become more confident about their reward estimates for players, the probability estimates that the players keep track of will be more representative of true probabilities. Given the structure of Algorithm 2 (PCA-SCA), we evaluate two methods to keep track of agents’ beliefs: UCB (PCA-UCB), and Thompson Sampling (PCA-TS). In this section, we detail the results of our experiments when using these approaches.
8.2.1. Using PCA-UCB
The results of the experiments using PCA-UCB in the APU model are summarized in Figure 2. We run PCA-UCB for 20,000 time steps for markets with random preferences on both sides, and for 10,000 time steps for when player preference heterogeneity is varied. The former results are in the first row of the figure, and the results of the latter are detailed in the second row of Figure 2.
Figure 2. Results of experiments run in the APU model using UCB to keep track of beliefs: first row has preferences uniformly random, second row has varied player preference heterogeneity with $N=K=10$. Key Takeaways: the markets converge to stability in expectation. If you compare this figure against Figure 3, Thompson Sampling converges to stability slightly faster than UCB and smoother, showing no dependence on player preference heterogeneity.
Firstly, notice that the APU model using PCA-UCB takes longer to stabilize compared to the APKP model using OCA-UCB, as seen in Figures 1 and 2. This is due to the fact that both players and arms need to learn their beliefs from scratch. Regardless, when market preferences are random, the trends in both figures are similar, with market stabilization probabilities approaching 1 and player regret tending to 0. This shows that PCA-UCB can find stable matches in a decentralized fashion with unknown preferences on both sides of the market when preferences are random.
The second row of sub-figures in Figure 2 shows what happens when the players’ preference heterogeneity is varied. We can see that there is no trend associated with $\beta$ concerning the algorithms’ convergence to stability. This is analogous to CA-UCB where a decrease in player preference heterogeneity does not have any effect on the convergence of the algorithm.
8.2.2. Using PCA-TS
Results for when PCA-TS was used in the APU model are shown in Figure 3. Like Figure 2 (the UCB results), the first row in Figure 3 is when the agent preferences are sampled uniformly at random, and the second row is where the players’ preference heterogeneity is varied.
Figure 3. Results of experiments run in the APU model using Thompson Sampling to keep track of beliefs. The first row shows market stability and average player regret when preferences are uniformly random on both sides of the market whereas the second row is for when player preference heterogeneity is varied. All the markets stabilize and player regret tends to 0. Performance was better than UCB in terms of convergence rate and smoothness.
Results from PCA-UCB (Figure 2) and PCA-TS (Figure 3) show similarities in market stability and player regret for all market sizes with uniformly random preferences, with larger markets stabilizing slower. We can also see that PCA-TS converges faster and more smoothly than PCA-UCB. We will discuss this particular property in further detail in the next subsection. The results demonstrate that Algorithm 2 can find a stable match in the APU model using both UCB and Thompson Sampling.
8.2.3. Convergence Comparison: UCB vs. Thompson Sampling
When looking at Figures 2 and 3, we can see that the Thompson Sampling approach of keeping track of beliefs appears better than UCB in terms of both speed and smoothness of convergence. To quantify this behavior, we turn to the convergence proxy we defined earlier, this time comparing the stability graphs of PCA-UCB and PCA-TS. We set $\mathcal{X}=1000,\theta=90$ and generate these lines, the results of which are summarized in Figure 4. Each color on the graph corresponds to the market size with the solid line representing PCA-UCB and the dotted line representing PCA-TS. We can see that the lines associated with Thompson Sampling reach 1.0 (i.e. 100% of the time-steps in question have stability $>90\%$) quicker and tend to stay that way.
Figure 4. Convergence Comparison of UCB and Thompson Sampling in the APU model with varying market sizes and uniformly random preferences on both sides of the market. The Thompson Sampling approach (dotted) converges slightly faster and stays that way compared to the UCB (solid) which tends to dip time and again before ultimately reaching convergence.
In comparison, the convergence proxy for UCB reaches 1.0 slower for each of the market sizes, and often the proxy value dips before going back up again. By nature of the way beliefs are kept track of, whenever a player chooses to explore, it causes a disturbance in the believed preference ordering of the player. When the internal belief state for a player changes, the player’s set of arms in consideration changes, leading to a different matching. This behavior causes the match to be unstable at some time steps. This is more frequent in the early stages of the algorithm and becomes less frequent as confidence increases about player preferences. PCA-TS exhibits this behavior only for $N=20$ but overall seems to reach 1.0 in a much smoother manner compared to PCA-UCB. The combination of these two reasons suggests that Thompson Sampling might be a better approach for keeping track of agent beliefs in the APU model.
9. Conclusion and Future Work
In this paper, we looked at the problem of two-sided matching with uncertain preferences. While recent work has developed algorithms that converge to a stable match when one side of the market must learn their preferences, we developed a new algorithm that converges to a stable match in two more challenging, settings: the APKP model where the arms have access to their preferences but the agents do not, and the APU model where none of the agents involved know their preferences a priori.
In the APKP model, we prove that the players can reliably learn their positioning in the arms’ preference orderings. This allows the players to construct a set similar to the plausible set from the CA-UCB algorithm and facilitates the convergence of the algorithm to a stable match. We also empirically studied the resulting algorithm, running simulation-based experiments and showing that the matching converges to stability, only slightly slower than the CA-UCB approach in a model where arm preferences are common knowledge. This slowness is expected, because the players now must learn more information before the algorithm can converge.
In the more complex situation represented by the APU model, we propose an algorithm where the players make use of the highest expected reward using an optimism function to pick the arms to attempt at each time step. The players do this by keeping track of estimates of the probability of winning a conflict against another player for each arm to pass to the optimism function in addition to the reward estimates. Then, the players attempt arms that maximize the product of the two. Furthermore, since the arms themselves do not know their preferences in this model, the arms keep track of reward estimates along with the confidence intervals over the players and pick the player with the highest estimated reward from the incoming requests, or if the intervals overlap then the arms pick overlapping players uniformly at random. We theoretically show that the algorithm converges to a stable match when using UCB as the reward estimate. Then we empirically show that the algorithm converges to a stable match, using both UCB and Thomson Sampling as belief estimators where we saw both these approaches converged to stability, with Thompson Sampling approaching stability quicker and more smoothly, mirroring the success of Thompson Sampling in simpler MAB settings.
Prior work has proved convergence in models where arm preferences are common knowledge. In this paper, we have provided new algorithms that provably convergence to stability in both the APKP model and the APU model – the most general setting of the two-sided matching problem.
Appendix A Full Proof of Convergence of PCA-UCB
In this section, we re-establish all the necessary information for the proof’s context. We consider the APU model with $K$ arms and $N$ players.
A.1. Preliminaries for the Arms
Each arm $k$ has a true mean reward associated with player $i$ denoted by $\mu_{i}^{k}$, and let $\hat{\mu}_{i}^{k}(t)$ denote the empirical mean computed from $n_{i}^{k}(t)$ samples by time $t$. Then, the corresponding confidence intervals that arm $k$ keeps track of are:
(3) $\ \begin{array}[]{rcl}\text{LCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)-c_{i}^{k}(t),\\ \text{UCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)+c_{i}^{k}(t)\end{array}$
where
$c_{i}^{k}(t)=\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}.$
Also, for any pair of players $p_{i}$ and $p_{j}$, define:
$\begin{array}[]{rcl}\Delta_{ij}^{k}&=&\mu_{i}^{k}-\mu_{j}^{k}\quad\text{for }\mu_{i}^{k}>\mu_{j}^{k},\\ \Delta_{\min}&=&\underset{i\neq j,\forall k}{\min}\lvert\mu_{i}^{k}-\mu_{j}^{k}\rvert,\\ n_{ij}^{k}(t)&=&\min\{n_{i}^{k}(t),n_{j}^{k}(t)\}.\end{array}$
Note that, based on our construction, $\Delta_{\min}=1$.
A.2. Preliminaries for the Players
Each player $p_{i}$ maintains an empirical win probability $\hat{P}_{ij}^{k}(t)$ against player $p_{j}$ for arm $a_{k}$, where the true probability $P_{ij}^{k}\in\{0,1\}$ is unknown. Let $N_{c_{ij}}^{k}(T)$ denote the number of conflicts between $p_{i}$ and $p_{j}$ for arm $a_{k}$ up to time $T$. We use the capital $N_{c}$ since our proof considers the period before the arms’ confidence intervals become disjoint (at time $T$), at which point the number of collisions is fixed.
A.2.1. Conflict Feedback to the Players
In our analysis, we assume without loss of generality that players are ordered so that $\mu_{i}^{k}>\mu_{j}^{k}$ when $p_{i}\succ_{a_{k}}p_{j}$. We say that the confidence intervals for arm $a_{k}$ are disjoint if, for every pair $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$,
$\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t).$
Conversely, if for some $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$ we have
$\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t),$
then the confidence intervals for those two players are said to be overlapping. When an arm’s confidence intervals overlap for a set of players, the conflict feedback is provided uniformly at random. Specifically, the arm selects the proposal with the highest UCB estimate; if there is any overlap with another player’s interval, it chooses uniformly at random from the contenders. As a result, the win probability estimates for losing players concentrate around $0.5$. Once the intervals become disjoint, however, the feedback becomes deterministic.
Definition 0 (Deterministic Feedback Event $\alpha(t)$).
For any time $t$, define
$\alpha(t)=\Bigl{\{}\forall\,k\in[K],\,\forall\,i,j\text{ with }\mu_{i}^{k}>\mu_{j}^{k}:\;\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t)\Bigr{\}}.$
Thus, $\alpha(t)$ is the event that, for every arm, the players’ confidence intervals are disjoint. When $\alpha(t)$ holds, the arm’s feedback is deterministic.
A.2.2. Arm Choice Mechanism of Players
Players use an optimism function to choose an arm at each time step. The function is defined as:
(1) $f(x)=\begin{cases}\displaystyle\frac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5,\end{cases}$
where $\kappa>0$ is the optimism parameter. By choosing $\kappa$ sufficiently large, even if a player’s raw empirical win estimate is as low as $0.5-\epsilon$ (for some $\epsilon>0$), the adjusted value $f(x)$ is boosted above a desired threshold, ensuring the arm continues to be selected. (This is formalized later in one of our lemmas.)
A.2.3. Plausible Sets and Inconsistent Triplets
In CA-UCB and OCA-UCB, for a player $p_{i}$, the plausible set at time-step $t$ is defined as
$S^{(i)}(t):=\{a_{j}:p_{i}\succ_{a_{j}}p_{k},\text{ where }\bar{A}^{k}(t-1)=a_{j}\}.$
That is, $S^{(i)}(t)$ is the set of all arms $a_{j}$ for which, at time-step $t-1$, some other player $p_{k}\neq p_{i}$ successfully pulled the arm, yet $p_{i}$ knows that the arm’s inherent preference favors $p_{i}$ over $p_{k}$. In CA-UCB, players had direct access to arms’ preferences, whereas in OCA-UCB, the arms’ preferences were learned deterministically. In the context of the APKP model, one may view an arm $a_{k}\in S^{(i)}(t)$ as having a win probability of 1 (with all other arms having a win probability of 0), thereby considering arms with a win probability of 1 as “plausible”. In the APU model, however, plausibility must be interpreted more generally as any arm with a non-zero probability of being won.
Definition 0 (Inconsistent Triplet).
We now define an inconsistent triplet in the APU model. Let $P_{ij}^{k}$ denote the true probability of player $p_{i}$ winning against player $p_{j}$ in a conflict for arm $a_{k}$, and let $\hat{P}_{ij}^{k}$ be the corresponding empirical probability. The triplet $(p_{i},p_{j},a_{k})$ is said to be inconsistent if
[$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
i.e., the player’s empirical belief about its win probability in the conflict deviates significantly (by at least $\epsilon$) from the true outcome—either overestimating a win or underestimating a loss. Next, we define a inconsequential triplet
Definition 0 (Inconsequential Triplet).
An inconsistent triplet is considered inconsequential if there exists an alternative arm that offers a better reward prospect. Formally, consider two triplets $(p_{i},p_{j},a_{k})$ and $(p_{i},p_{j}^{\prime},a_{k}^{\prime})$. If, at time $t$, the triplet $(p_{i},p_{j},a_{k})$ is inconsistent but $\text{UCB}_{i}^{k^{\prime}}>\text{UCB}_{i}^{k}$ and the triplet $(p_{i},p_{j}^{\prime},a_{k}^{\prime})$ is consistent (i.e., arm $a_{k^{\prime}}$ is a better prospect for $p_{i}$), then player $p_{i}$ will select $a_{k^{\prime}}$. In this case, the inconsistency in $(p_{i},p_{j},a_{k})$ is rendered inconsequential.
And finally a triplet is considered consequential if it is not inconsequential.
A.3. Main Results
The main proof can be summarized in the following main theorems:
[…]
A.4. Proof of Theorem 6.1
We will begin the proof of Theorem 6.1 by first establishing some important lemmas that relate to bounding the time that is taken for the arms to learn their preferences (i.e. the event $\alpha(t)$ happening). We begin by looking at the tracked reward estimates by an arm $a_{k}$ for a single pair of players $p_{i}$ and $p_{j}$.
Lemma A.0 (Single Pair of Players Overlap).
Let $p_{i}$ and $p_{j}$ be two players such that $\mu_{i}^{k}>\mu_{j}^{k}$ and $\hat{\mu}_{i}^{k}$ and $\hat{\mu}_{j}^{k}$ be the empirical reward estimate by arm $a_{k}$ for the players respectively at time $t$. With the confidence intervals are defined as in Equation 3, the probability that the confidence intervals for $p_{i}$ and $p_{i}$ overlap is bounded by:
$P\left(\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t)\right)\leq\exp\left(-\frac{\left(\Delta_{ij}^{k}-(c_{i}^{k}(t)+c_{j}^{k}(t))\right)^{2}}{2\left(\frac{1}{n_{i}^{k}(t)}+\frac{1}{n_{j}^{k}(t)}\right)}\right)$
where $\Delta_{ij}^{k}=\mu_{i}^{k}-\mu_{j}^{k}$ is the true mean difference between players $p_{i}$ and $p_{j}$, and $n_{i}^{k}(t)$ and $n_{j}^{k}(t)$ are the numbers of samples for each player at time $t$.
Proof.
In the context of this proof, it is clear that we are talking about arm $a_{k}$, so we will not use the superscript of $k$ for clarity. The event $\text{LCB}_{i}(t)\leq\text{UCB}_{j}(t)$ can be rewritten as:
$\hat{\mu}_{i}(t)-\hat{\mu}_{j}(t)\leq c_{i}(t)+c_{j}(t)$
We have that $\Delta_{ij}=\mu_{i}-\mu_{j}$, which is positive since $\mu_{i}>\mu_{j}$. Then we can express the difference in sample means as:
$\hat{\mu}_{i}(t)-\hat{\mu}_{j}(t)=\Delta_{ij}+\left(\hat{\mu}_{i}(t)-\mu_{i}\right)-\left(\hat{\mu}_{j}(t)-\mu_{j}\right)$
Thus, the event $\text{LCB}_{i}(t)\leq\text{UCB}_{j}(t)$ becomes:
$\left(\hat{\mu}_{i}(t)-\mu_{i}\right)-\left(\hat{\mu}_{j}(t)-\mu_{j}\right)\leq c_{i}(t)+c_{j}(t)-\Delta_{ij}$
Define $\epsilon_{ij}=\Delta_{ij}-(c_{i}(t)+c_{j}(t))$. Then we need to bind the probability of the event
$Z=\left(\hat{\mu}_{i}(t)-\mu_{i}\right)-\left(\hat{\mu}_{j}(t)-\mu_{j}\right)\leq-\epsilon_{ij}$
We will use the Hoeffding Inequality for Sub-Gaussian random variables to do this. But first, we need to show that $Z$ is Sub-Gaussian. We do this in the following lemma.
Lemma A.0.1.
Let $\{X_{i,k}\}_{k=1}^{n_{i}(t)}$ be independent samples drawn from $\mathcal{N}(\mu_{i},1)$ and $\{X_{j,k}\}_{k=1}^{n_{j}(t)}$ be independent samples drawn from $\mathcal{N}(\mu_{j},1)$. Define the sample means
$\hat{\mu}_{i}(t)=\frac{1}{n_{i}(t)}\sum_{k=1}^{n_{i}(t)}X_{i,k}\quad\text{and}\quad\hat{\mu}_{j}(t)=\frac{1}{n_{j}(t)}\sum_{k=1}^{n_{j}(t)}X_{j,k},$
and consider the random variable
$Z=\Bigl{(}\hat{\mu}_{i}(t)-\mu_{i}\Bigr{)}-\Bigl{(}\hat{\mu}_{j}(t)-\mu_{j}\Bigr{)}.$
Then, $Z$ is a sub-Gaussian random variable with a sub-Gaussian parameter
$\sigma_{Z}=\sqrt{\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}},$
Proof.
By properties of the normal distribution, we have:
$\hat{\mu}_{i}(t)\sim\mathcal{N}\!\left(\mu_{i},\frac{1}{n_{i}(t)}\right)\quad\text{and}\quad\hat{\mu}_{j}(t)\sim\mathcal{N}\!\left(\mu_{j},\frac{1}{n_{j}(t)}\right).$
And, we have that $Z$ is a random variable with:
$Z=\Bigl{(}\hat{\mu}_{i}(t)-\mu_{i}\Bigr{)}-\Bigl{(}\hat{\mu}_{j}(t)-\mu_{j}\Bigr{)}.$
Since,
$\hat{\mu}_{i}(t)-\mu_{i}\sim\mathcal{N}\!\left(0,\frac{1}{n_{i}(t)}\right)\quad\text{and}\quad\hat{\mu}_{j}(t)-\mu_{j}\sim\mathcal{N}\!\left(0,\frac{1}{n_{j}(t)}\right)$
are independent, it follows that:
$Z\sim\mathcal{N}\!\left(0,\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}\right).$
Let,
$\sigma_{Z}^{2}=\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)},$
so that
$\sigma_{Z}=\sqrt{\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}}.$
Recall that a random variable $X$ is said to be sub-Gaussian with parameter $\sigma$ if
$\mathbb{E}\!\left[e^{t(X-\mathbb{E}[X])}\right]\leq e^{\frac{\sigma^{2}t^{2}}{2}}\quad\text{for all }t\in\mathbb{R}.$
A well-known fact is that if $X\sim\mathcal{N}(0,\sigma^{2})$, then
$\mathbb{E}\!\left[e^{tX}\right]=e^{\frac{\sigma^{2}t^{2}}{2}}\quad\text{for all }t\in\mathbb{R},$
which shows that $X$ is sub-Gaussian with parameter $\sigma$. Since $Z\sim\mathcal{N}(0,\sigma_{Z}^{2})$, we conclude that $Z$ is sub-Gaussian with parameter
$\sigma_{Z}=\sqrt{\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}},$
Lemma A.0.1 establishes $Z$ as a sub-Gaussian random variable. Then, using the Hoeffding Inequality for Sub-Gaussian Random Variables yields:
$P(Z\leq-\epsilon)\leq\exp\left(-\frac{\epsilon^{2}}{2\sigma^{2}}\right)$
Applying this to our event, we have
$P\left(\text{LCB}_{i}(t)\leq\text{UCB}_{j}(t)\right)\leq\exp\left(-\frac{\epsilon_{ij}^{2}}{2\sigma_{ij}^{2}}\right)$
Since $\epsilon_{ij}=\Delta_{ij}-(c_{i}(t)+c_{j}(t))$ and $\sigma_{ij}^{2}=\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}$, we obtain
$P\left(\text{LCB}_{i}(t)\leq\text{UCB}_{j}(t)\right)\leq\exp\left(-\frac{\left(\Delta_{ij}-(c_{i}(t)+c_{j}(t))\right)^{2}}{2\left(\frac{1}{n_{i}(t)}+\frac{1}{n_{j}(t)}\right)}\right)$
Now that we have established the probability bound for any given pair of players for an arm, we will extend this argument to all arms keeping track of the confidence intervals of all players.
Lemma A.0 (All Pair of Players Overlap).
[…]
$P\left(\bigcup_{i\neq j,\mu_{i}^{k}>\mu_{j}^{k}}\left\{\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t)\right\}\right)\leq\frac{N^{2}}{2}\cdot\exp\left(-\frac{n^{k}\left(\Delta_{\text{min}}^{k}-2c^{k}(t)\right)^{2}}{4}\right)$
Proof.
In Lemma A.4, we have established that for an arm $a_{k}$ and any two players $p_{i}$ and $p_{j}$ s.t. $\mu_{i}^{k}>\mu_{j}^{k}$:
$P\left(\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t)\right)\leq\exp\left(-\frac{\left(\Delta_{ij}^{k}-(c_{i}^{k}(t)+c_{j}^{k}(t))\right)^{2}}{2\left(\frac{1}{n_{i}^{k}(t)}+\frac{1}{n_{j}^{k}(t)}\right)}\right)$
[…]
  • $\Delta_{ij}^{k}\geq\Delta_{\min}^{k}$,
  • $(c_{i}^{k}(t)+c_{j}^{k}(t))\leq 2c^{k}(t)$ where $c^{k}(t)=\sqrt{\frac{3\ln(t)}{2n^{k}(t)}}$, and
  • $\left(\frac{1}{n_{i}^{k}(t)}+\frac{1}{n_{j}^{k}(t)}\right)\leq\frac{2}{n^{k}(t)}$
The following inequality still holds:
$P\left(\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t)\right)\leq\exp\left(-\frac{n^{k}(t)\left(\Delta_{\min}^{k}-2c^{k}(t)\right)^{2}}{4}\right)$
Next, the total number of players where $\mu_{i}>\mu_{j}$ is at most $\frac{N(N-1)}{2}$. For each arm, we have $N$ choices for player $i$ and $N$ choices for player $j$, giving a total of $N\times N=N^{2}$ ordered pairs $(i,j)$ where $i$ and $j$ are any two players (including the cases where $i=j$). Since we are interested only in pairs $(i,j)$ where $i\neq j$, we exclude the $N$ pairs where $i=j$ (e.g., $(1,1),(2,2),\dots,(N,N)$). Thus, there are $N^{2}-N=N(N-1)$ pairs with $i\neq j$. For each distinct pair $(i,j)$ where $i\neq j$, only one of the following statements can be true: either $\mu_{i}>\mu_{j}$ or $\mu_{j}>\mu_{i}$. Therefore, we only consider one direction for each pair (e.g., $(i,j)$ and not $(j,i)$), effectively halving the number of pairs. Since we only count each distinct pair $(i,j)$ once, the number of pairs with $i\neq j$ is halved. This gives us $\frac{N(N-1)}{2}$ pairs where we consider $\mu_{i}>\mu_{j}$ for each distinct player pair $(i,j)$.
Finally, with a straightforward application of the union bound over all possible pairs, we get:
$P\left(\bigcup_{i\neq j,\mu_{i}^{k}>\mu_{j}^{k}}\left\{\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t)\right\}\right)$ $\leq\frac{N(N-1)}{2}\cdot\exp\left(-\frac{n^{k}\left(\Delta_{\text{min}}^{k}-2c^{k}(t)\right)^{2}}{4}\right)$
$\leq\frac{N^{2}}{2}\cdot\exp\left(-\frac{n^{k}\left(\Delta_{\text{min}}^{k}-2c^{k}(t)\right)^{2}}{4}\right)$
Lemma A.0 (All Arms for All Pair of Players Overlap).
For $K$ arms with independent confidence interval calculations, the probability of any overlap on any arm satisfies:
$P\left(\exists a_{k}\text{ with }\mu_{i}^{k}>\mu_{j}^{k}\text{ and }\text{LCB}_{i}^{k}(t)\leq UCB_{j}^{k}(t)\right)\leq K\cdot\frac{N^{2}}{2}\cdot\exp\left(-\frac{n(t)\left(\Delta_{\text{min}}-2c(t)\right)^{2}}{4}\right)$
where $n(t)=\underset{k\in K}{\min}\left\{n^{k}(t)\right\}$, $\Delta_{\min}=\underset{k\in K}{\min}\left\{\Delta_{\min}^{k}\right\}$ and $c(t)=\sqrt{\frac{3\ln(t)}{2n(t)}}$
Proof.
The proof of this lemma is a straightforward application of Lemma A.5 using the union bound over all $K$ arms and recognizing that replacing $n(t),\Delta_{\min},\text{ and }c(t)$ in the expression still maintains the upper-bound on the probability. ∎
A.4.1. Putting it together
Now, we will use Lemma A.6 to prove Theorem 6.1. Notice that the probability that we have upper-bounded in Lemma A.6 is the complement of the event $\alpha(t)$ (i.e. the confidence intervals for all the players tracked by all the arms are disjoint). Hence:
$P\left(\alpha(t)\text{ occurs at time }t\right)$ $=1-P\left(\exists a_{k}\text{ with }\mu_{i}^{k}>\mu_{j}^{k}\text{and }\text{LCB}_{i}^{k}(t)\leq UCB_{j}^{k}(t)\right)$
$\leq 1-\underbrace{K\cdot\frac{N^{2}}{2}\cdot\exp\left(-\frac{n(t)\left(\Delta_{\text{min}}-2c(t)\right)^{2}}{4}\right)}_{\delta}$
Important: Notice that as $t\to\infty$, $n(t)\to\infty$. Since the contents inside the brackets are a positive number by being squared, the overall exponent $\to-\infty\implies\delta\to 0$. This will be important later in establishing that PCA-SCA’s execution is identical to CA-UCB as $t\to\infty$.
A.5. Proof of Theorem 6.1
Assume that the event $\alpha$ occurs at time $T$, so that the conditions of Theorem 6.1 are satisfied. Our goal is to analyze the evolution of the players’ empirical winning probabilities, $\hat{P}_{ij}^{k}$, both before and after $\alpha(T)$.
We begin by considering $\hat{P}_{ij}^{k}$ prior to $\alpha(T)$. This is formalized in the following lemma:
Lemma A.0.
Assume event $\alpha$ has happened by time $T$. The empirical win estimate tracked by a player $p_{i}$ against player $p_{j}$ for an arm $a_{k}$ at time $T$ satisfies:
$P(\hat{P}_{ij}^{k}\leq 0.5-\epsilon)\leq exp\left(-2N_{c_{ij}}^{k}(T)\epsilon^{2}\right)$
for some $\epsilon\geq 0$ and where $N_{c_{ij}}^{k}(T)$ is the the number of times players $p_{i}$ and $p_{j}$ have conflicted for arm $a_{k}$ up to time $T$.
Proof.
The proof is a standard application of concentration of sum of i.i.d Bernoulli random samples. We know by the construction of the algorithm, while there is still overlap between the confidence interval of the arm-tracked reward for the players (i.e. until $\alpha$ has happened), the arm picks the players uniformly at random from the overlapping players. Then:
$\hat{P}_{ij}^{k}=\frac{1}{N_{c_{ij}}^{k}(T)}\sum_{i=1}^{N_{c_{ij}}^{k}(T)}x_{i}\left(\text{ where }x_{i}\sim\text{Bernoulli}(0.5)\right)$
By standard concentration, we have that for any $\epsilon>0$:
$P(\hat{P}_{ij}^{k}\leq 0.5-\epsilon)\leq exp\left(-2N_{c_{ij}}^{k}(T)\epsilon^{2}\right)$
Lemma A.7 quantifies how unlikely it is for a player to have an empirical win estimate significantly below 0.5, which is crucial to prevent a player from discarding an arm due to an erroneously low conflict win estimate.
To enforce this safeguard, we employ the optimism function defined in Equation 1, yielding the next lemma:
Lemma A.0 (Optimistic Picking of Arm by Player).
Let event $\alpha$ occur at time $T$. Suppose player $p_{i}$ has an empirical win estimate $\hat{P}_{ij}^{k}$ for winning a conflict against player $p_{j}$ on arm $a_{k}$. Define the optimism function $f\colon[0,1]\to[0,1]$ by
$f(x)=\begin{cases}\displaystyle\frac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&\text{if }0\leq x\leq 0.5,\\[4.30554pt] 1,&\text{if }x>0.5.\end{cases}$
Then for any $\epsilon>0$, there exists a sufficiently large constant $\kappa$ such that when $p_{i}$ uses $f\bigl{(}\hat{P}_{ij}^{k}\bigr{)}$ to select an arm, it will choose arm $a_{k}$ with probability at least
$1-\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)},$
where $N_{c_{ij}}^{k}(T)$ denotes the number of conflicts between $p_{i}$ and $p_{j}$ up to time $T$.
Proof.
Let us pick a threshold $\epsilon>0$ such that even if $\hat{P}_{ij}^{k}=0.5-\epsilon$, we have that $f(0.5-\epsilon)\geq q>0$, ensuring that the players’ win heuristic remains sufficiently large to keep attempting the arm:
$f(0.5-\epsilon)=\frac{1-e^{-\kappa(0.5-\epsilon)}}{1-e^{-\kappa/2}}\geq q$
This inequality can be solved for $\kappa$ and any desired value of $\epsilon$ and $q$. Suppose then we pick a suitable $\kappa$, the player only fails to try the arm if $\hat{P}_{ij}^{k}<0.5-\epsilon$. By Lemma A.7 we know that $P(\hat{P}_{ij}^{k}<0.5-\epsilon)\leq\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$. Therefore, with probability at least $1-\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}$, the player’s win estimate will be at least $0.5-\epsilon$, ensuring that the player continues to attempt the arm. ∎
Important: Notice that, regardless of what value of $\epsilon$ we pick, as $\kappa\to\infty,f(0.5-\epsilon)\to 1$ i.e. regardless of where we pick the threshold or how far the empirical estimate is from 0.5, we can always pick $\kappa$ to be arbitrarily large to reach a level to optimism required that the players attempt arms.
We then extend this result to all possible triplets to ensure that no player stops prematurely due to adverse fluctuations:
Lemma A.0.
Assume that event $\alpha$ occurs at time $T$ and that all players employ the optimism function when selecting arms. Define
$N_{c}(T)=\min_{\begin{subarray}{c}i,j\in N\\ k\in K\end{subarray}}\{N_{c_{ij}}^{k}(T)\},$
the minimum number of conflicts observed among any pair of players on any arm by time $T$. Then, for any $\epsilon>0$, the probability that all players continue attempting arms is at least
$1-\frac{KN^{2}}{2}\exp\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}.$
Proof.
By Lemma A.7, for any players $p_{i}$ and $p_{j}$ and any arm $a_{k}$ (with $p_{i}\succ_{a_{k}}p_{j}$), we have
$P\Bigl{(}\hat{P}_{ij}^{k}\leq 0.5-\epsilon\Bigr{)}\leq\exp\Bigl{(}-2\,N_{c_{ij}}^{k}(T)\,\epsilon^{2}\Bigr{)}.$
Since $N_{c_{ij}}^{k}(T)\geq N_{c}(T)$ for all such triplets, it follows that
$P\Bigl{(}\hat{P}_{ij}^{k}\leq 0.5-\epsilon\Bigr{)}\leq\exp\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}.$
There are at most $K\cdot\frac{N(N-1)}{2}$ triplets in consideration in the system. Applying the union bound, the probability that any triplet has an empirical win estimate below $0.5-\epsilon$ is at most
$\frac{KN^{2}}{2}\exp\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}.$
Finally, by Lemma A.8, as long as every empirical win estimate exceeds $0.5-\epsilon$, the optimism function guarantees that players will continue selecting arms. Thus, the probability that all players continue attempting arms is at least
$1-\frac{KN^{2}}{2}\exp\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}.$
These lemmas establish a lower bound on the probability that players continue selecting the arms after $\alpha(T)$. Next, we set a suitable $\epsilon$ so that the tail probabilities align. Let $B$ denote the event that no triplet has an empirical win estimate $\hat{P}$ falling below $0.5-\epsilon$ after $\alpha(T)$. We require that
$P(B)\geq P(\alpha),$
$P(\alpha)\geq 1-K\cdot\frac{N^{2}}{2}\exp\!\left(-\frac{n(T)\left(\Delta_{\text{min}}-2c(T)\right)^{2}}{4}\right)$
$P(B)\geq 1-\frac{KN^{2}}{2}\exp\!\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}.$
Matching the tail probabilities yields
$\exp\!\left(-\frac{n(T)\left(\Delta_{\text{min}}-2c(T)\right)^{2}}{4}\right)\approx\exp\!\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)},$
which implies
$\epsilon\approx\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}.$
By the algorithm’s design, once $\alpha$ holds at time $T$ (so all confidence intervals are disjoint), any further conflict feedback becomes deterministic. As long as players keep attempting arms, their empirical win estimates will converge to the true values. The following lemma establishes this.
Lemma A.0.
Let $T\geq 1$ be a positive integer and let $\hat{x}_{T}\in[0,1]$ be a real number, which we interpret as the average of $T$ samples. Let $p\in\{0,1\}$ be a fixed value (the “true” value). For any integer $t\geq 0$, define the updated average after adding $t$ additional samples (each equal to $p$) by
$\hat{x}_{T+t}\;=\;\frac{T\,\hat{x}_{T}+t\,p}{T+t}\,.$
Then, the error after $t$ additional samples satisfies
$\left|\hat{x}_{T+t}-p\right|\;\leq\;\frac{T}{T+t}$
Proof.
Starting from the definition of $\hat{x}(T+t)$, we have
$\hat{x}_{T+t}-p$ $=\frac{T\,\hat{x}_{T}+t\,p}{T+t}-p$
$=\frac{T\,\hat{x}_{T}+t\,p-p(T+t)}{T+t}$
$=\frac{T\,\hat{x}_{T}+t\,p-pT-pt}{T+t}$
$=\frac{T\,(\hat{x}_{T}-p)}{T+t}$
Taking absolute values of both sides gives
$\left|\hat{x}(T+t)-p\right|$ $=\frac{T}{T+t}\,\left|\hat{x}-p\right|$
$\leq\frac{T}{T+t}$
But realize that not all conflict win estimates will converge since some triplets eventually stop receiving feedback. In the context of Theorem 6.1, before $\alpha(T)$, players are getting sample rewards as they are getting picked by the arms uniformly at random, leading to the players having some optimistic reward estimate. Consequently, after $\alpha(T)$, players have some reward estimates for the arms and a belief about conflict wins (with the worst-case estimate being around $0.5-\epsilon$). Right after $\alpha(T)$ occurs, the combined effect of the optimistic reward estimates and optimism function ensures that conflicts persist i.e. the players continue trying the arms. As more conflicts occur, the empirical win estimates gradually converge toward their true values, causing the impact of the optimism function to either be magnified (if the win probability exceeds 0.5, effectively guaranteeing a win) or diminished. This process enables players to gather more accurate information for selecting arms. In cases where a player already receives a better reward from one arm, they may cease exploring some inferior arm. This would cause the conflict win probability of the associated inferior triplet to never reach the true value. But as we will see in the following lemma, this is the case only for inconsequential triplets.
Lemma A.0.
Suppose the $\alpha(T)$ condition is achieved. Then an inconsistent triplet $(p_{i},p_{j},a_{k})$ at time $T$ is either resolved, or is inconsequential by time $(T+n)$.
Proof.
Recall that an inconsistent triplet $(p_{i},p_{j},a_{k})$ at time $T$ means one of the following holds for some $\epsilon>0$:
  1. (1)
    $P_{ij}^{k}=1$ but $\hat{P}_{ij}^{k}\leq 1-\epsilon$;
  2. (2)
    $P_{ij}^{k}=0$ but $\hat{P}_{ij}^{k}\geq\epsilon$.
In other words, player $p_{i}$’s empirical estimate $\hat{P}_{ij}^{k}$ is “far” from the true probability $P_{ij}^{k}\in\{0,1\}$. Because the event $\alpha(T)$ has occurred, all subsequent feedback from time $t$ onward is deterministic and aligns with the true outcome. Since we condition on $\alpha(T)$, we preclude $p_{i}$ from not attempting $a_{k}$. We split into two cases, depending on whether the triplet $(p_{i},p_{j},a_{k})$ remains relevant to $p_{i}$’s decisions:
Case 1: The triplet remains relevant.
If player $p_{i}$ continues to attempt arm $a_{k}$ and thus periodically conflicts with player $p_{j}$, then each time this conflict occurs, the empirical estimate $\hat{P}_{ij}^{k}$ receives noise-free feedback. Consequently:
  • If $P_{ij}^{k}=1$, $p_{i}$ always wins these conflicts. Hence, $\hat{P}_{ij}^{k}$ will rapidly increase above $1-\epsilon$, thus eliminating the inconsistency.
  • If $P_{ij}^{k}=0$, $p_{i}$ always loses. Hence, $\hat{P}_{ij}^{k}$ will drop below $\epsilon$, again removing the inconsistency.
By Lemma A.10, only a bounded number of outcomes is required before $\hat{P}_{ij}^{k}$ becomes accurate enough that $(p_{i},p_{j},a_{k})$ is no longer inconsistent.
Case 2: The triplet becomes inconsequential.
If instead player $p_{i}$ no longer selects arm $a_{k}$—for instance, because it discovers another arm $\smash{a_{k^{\prime}}}$ with a higher UCB estimate or learns that $a_{k}$ is not profitable—then the conflict $(p_{i},p_{j},a_{k})$ does not occur in future rounds. In that case, the belief about $P_{ij}^{k}$ no longer influences $p_{i}$’s decisions and hence no longer impacts $p_{i}$’s rewards. We then say that $(p_{i},p_{j},a_{k})$ is inconsequential, because its incorrect empirical probability does not affect the algorithm’s action selection.
In either scenario, by at most $n$ additional time steps, any given inconsistent triplet is either corrected by deterministic feedback (Case 1) or becomes irrelevant to the player (Case 2). This shows that every inconsistent triplet at time $t$ is resolved or inconsequential by time $t+n$. ∎
Now, we are ready to complete the proof of Theorem 6.1. By design, once $\alpha$ occurs at time $T$, the feedback provided to the players is deterministic. Moreover, by Lemma A.9, the probability that the optimism function lets the players continue attempting the arms despite bad conflict win estimates is lower bounded by:
$1-\underbrace{\frac{KN^{2}}{2}\exp\!\Bigl{(}-2\,N_{c}(T)\,\epsilon^{2}\Bigr{)}}_{\tau(t)},$
where $\epsilon=\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}$ is chosen to match the tail probability in Theorem 6.1. Finally, by applying Lemma A.10, we conclude that for all consequential conflicts $(p_{i},p_{j},p_{k})$ and any $t\geq 0$, the updated empirical probability satisfies
$\left|\hat{P}_{ij}^{k}(T+t)-p_{ij}^{k}\right|\leq\frac{T}{T+t}.$
And by Lemma A.11, conflict estimates that fail to satisfy this inequality are inconsequential. This completes the proof of Theorem 6.1. ∎
Important: Notice that as $t\to\infty,T\to\infty\implies n(T)\to\infty$. Because we condition on $\alpha(T)$, $N_{c}(T)$ is a constant. This implies, for a similar reason to $\delta(t)\to 0$, as $t\to\infty,\tau(t)\to 0$. This will later be important when establishing that PCA-SCA’s algorithm will eventually be identical to CA-UCB as $t\to\infty$
A.6. Proof of Theorem 6.1
The proof follows by combining Theorem 6.1, Theorem 6.1, and the supporting lemmas to show that, as $t\to\infty$, the execution of PCA-SCA converges to that of CA-UCB.
Theorem 6.1 provides a probabilistic bound on the event $\alpha(T)^{c}$ (i.e. when the arms’ confidence intervals overlap and the win feedback remains non-deterministic). Theorem 6.1 further bounds the event that the optimism function fails to motivate continued sampling, ensuring that every consequential triplet eventually receives sufficient feedback. Formally, define the failure event as things “going wrong” as:
[…]
$\leq P(\text{ arms do not learn preferences })+$
$P(\text{ some triplet has a win estimate less than }0.5-\epsilon)$
[…]
where both $\delta(t)$ and $\tau(t)$ tend to zero as $t\to\infty$. Consequently, $P(\text{failure})\to 0$. Moreover, by Lemma A.11, for every consequential triplet the empirical conflict win probability converges arbitrarily close to its true value as $t\to\infty$. This implies that, in the limit, arms’ have knowledge of their own preferences, and players have knowledge of arms’ preferences (in terms of the win conflict probabilities). If we consider all the arms where the probability of winning a conflict is $\geq 1-\epsilon$ as the arms that go in the player’s plausible set, players always choose the arm with the highest UCB estimate. It then follows that, with probability tending to one, the execution of PCA-SCA becomes identical to that of CA-UCB. ∎
Appendix B Alternative Methods for Controlling Player Preference Heterogeneity
There are alternative methods for controlling heterogeneity in player preferences, for example notions like uncoordinated/coordinated markets (Ackermann et al., 2008). First, note that uncoordinated markets as defined in (Ackermann et al., 2008) essentially have uncorrelated preferences (equivalent to our model with low $\beta$), and coordinated markets introduce correlation, albeit using a different model than ours (with correlation defined on the edges of the matching graph rather than the vertices).
We ran experiments (with results averaged over 100 different runs for N=K=10) with edge-correlated preferences as defined in (Ackermann et al., 2008), the results of which are shown in Figure 5.
Figure 5. Convergence proxy for players with correlated preferences running OCA-UCB in the APKP model with $\mathcal{X}=50,\theta=90$ The line in purple is the the edge-correlation as defined in (Ackermann et al., 2008) and the rest correspond to our paper. We can see that the edge-correlated cases are comparable to when preferences are uniformly random (compare with node-correlation, where convergence time increases)
Our findings show that the convergence time in these edge-correlated cases is actually comparable to when preferences are uniformly random (compare with node-correlation, where convergence time increases).
Appendix C Visualization of the Optimism Function
Figure 6. Visualization of different levels of “optimism” for different values of $\kappa$. Note that as $\kappa$ increases, the players are more optimistic about their chances of winning a conflict if if their empirical estimate is below 0.5